OpenStudy (anonymous):

BEST ANSWER REWARDED:WRITE THE EQUATION OF THE SIN FUNCTION :AMPLITUDE 3 , PERIOD PI/4

OpenStudy (anonymous):

amplitude is 3, put a 3 out front

OpenStudy (anonymous):

so something like \[3\sin(bx)\]

OpenStudy (anonymous):

to find \(b\), the period of \(\sin(bx)\) is always \(\frac{2\pi}{b}\) set \[\frac{2\pi}{b}=\frac{\pi}{4}\] and solve for \(b\)

OpenStudy (anonymous):

ok and how do u solve for b? cross multiply? can you demonstrate @satellite73